OpenStudy (anonymous):

What makes Equation 1 where Equation 2? Or what property? Equation 1 - b = x/a Equation 2 - x = b/a

OpenStudy (anonymous):

The "-" are not minus signs

OpenStudy (anonymous):

That is not true. x = ba is the conclusion from equation 1

OpenStudy (anonymous):

I might've written it wrong, Equation 1 is 8 = 16/2 , Equation 2 is 16= 8*2

OpenStudy (anonymous):

yes that is better

OpenStudy (anonymous):

just multiply both sides by 2. Multiplicative property of equality

OpenStudy (anonymous):

Thanks